# Dissipative solutions to Stochastic 3D Euler equations

### v2

@@ Abstract.
We construct probabilistically strong solutions to the three-dimensional Euler equations perturbed by additive noise that are $\mathbb{P}$-almost surely continuous in time, Hölder in space, and satisfy the local energy inequality up to an arbitrarily large stopping time.
@@
Stochastic Euler equations, Local energy inequality, Convex integration
@@ 1. Introduction
In this work, we consider the stochastic Euler equations on the three-dimensional torus $\mathbb{T}^{3}:=\mathbb{R}^{3}/2\pi\mathbb{Z}^{3}$ driven by additive noise
@@
Euler equations govern the evolution of an ideal fluid and constitute one of the most important models in fluid dynamics. The presence of a random external force in (SE) can help represent many phenomena perturbing the dynamics that are neglected by the deterministic equation, such as endogenous microscopic thermal effects or exogenous influences of the environment on the fluid. First investigations on (SE) date back to the works [MiVa00, Ki09], where local existence of smooth solutions is proved (see also [BeFl99] for other results in dimension two). Measure-valued solutions satisfying weak-strong uniqueness and a version of the energy inequality have been constructed in [BrMo21], building upon results by DiPerna and Majda [DPMa87] for the deterministic Euler equations $Q=0$. Whether solutions are unique or not remained an open problem until the work [HoZhZh22] by Hofmanová, Zhu, and Zhu, adapting for the first time the convex integration scheme developed by De Lellis and Székelyhidi Jr. in [DLSz09] to the stochastic case, and obtaining non-uniqueness in law for (SE). After [HoZhZh22] made clear that convex integration techniques can be applied to stochastic equations (let us also mention [ChFeFl21] in this regard), a sequence of papers has appeared [HoZhZh25, LuZh24, KiKo24, LuLuZh25] extending the works [DLS13, DLS14, BuDLIsSz15, BuDLIsSz16, Is18, BuDLSzVi19] on the deterministic Euler equations. In particular, for every $\vartheta\in(0,1/3)$, Lü, Lü, and Zhu proved in [LuLuZh25] existence and non-uniqueness of solutions of (SE) with regularity $u\in C([0,\infty),C^{\vartheta}(\mathbb{T}^{3},\mathbb{R}^{3})$ almost surely, which are strictly dissipative up to an arbitrary large stopping time $\mathfrak{t}$ in the sense that $u$ satisfies almost surely the energy inequality
@@
The importance of a local form of energy inequality towards partial regularity results for suitable weak solutions of the Navier-Stokes equations is known since the works of Scheffer [Sc77] and Caffarelli, Kohn, and Niremberg [CaKoNi82]. Similar results are available for the stochastic Navier-Stokes equations, see [Ro10, FlRo02]. The local energy balance (LEI) is a natural condition to impose also on the stochastic Euler equations (SE), since any physically relevant $L^{3}_{t,x}$ zero-viscosity limit of suitable solutions of the stochastic Navier-Stokes equations must inherit this property. It describes the local change in kinetic energy due to spatial roughness of $u$: indeed, the equality in (LEI) holds for smooth solutions $u$, and in general the inequality corresponds to no local creation of kinetic energy due to irregularities of $u$. We refer to the celebrated work of Duchon and Robert [DuRo00] for more details.
In the deterministic setting $Q=0$, existence and non-uniqueness of solutions to the system (SE)-(LEI) have been shown in the series of papers [DLSz10, Is22, DLK23, GiKwNo]. The purpose of the present work is to provide similar results in the stochastic case $Q\neq 0$. More specifically, we will prove the following:
@@
Let $(\Omega,\mathbb{F},\{\mathbb{F}_{t}\}_{t\geq 0},\mathbb{P})$ be a stochastic basis with right-continuous and complete filtration supporting a sufficiently smooth-in-space $Q$-Wiener process $Q^{1/2}W$ with zero space average and null divergence. Let be given $T>0$, $\varkappa\in(0,1)$, $\alpha\in(0,1/7)$ and $E\in C^{1}_{loc}([0,\infty),\mathbb{R})$. Then there exist two $\{\mathbb{F}_{t}\}_{t\geq 0}$-progressively measurable stochastic processes
(1.2) $u:\Omega\to C([0,\infty),C^{\alpha}(\mathbb{T}^{3},\mathbb{R}^{3})),\qquad p:\Omega\to C([0,\infty),C^{2\alpha}(\mathbb{T}^{3},\mathbb{R})),$
@@ Acknowledgements
UP has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 949981).
@@
2. Preliminaries and main iterative proposition
2.1. Reformulation of the problem without local martingales
@@
(2.4) $\frac{d}{dt}\frac{|v|^{2}}{2}+v\cdot\nabla p+v\cdot\mathord{{\rm div}}\big((v+z)\otimes(v+z)\big)=-E^{\prime}.$
@@
2.2. Notation
Let $N\in\mathbb{N}$ and $\delta\in(0,1)$, and let us define the following Hölder seminorms:
Then, the Hölder norms are defined as
Furthermore, in order to simplify the statement of many propositions here contained, we adopt the notational convention $\|f\|_{M}\equiv 0$ whenever the index $M<0$.
2.3. Preparation of the convex integration scheme
The proof of ˜1.1 relies on a convex integration scheme, mostly inspired by the works [Is22, DLK23, LuLuZh25]. Let us introduce various parameters that will allow to estimate all the objects appearing in the upcoming sections. Define for $q\in\mathbb{N}$:
where $a>2$ will be very large, $b>1$ will be chosen close to 1 and $\alpha\in(0,1/7)$ indicates the Hölder regularity of the solution identified by the iterative procedure.
Let also also denote $N_{1}=n_{0}+5$, where the parameter $n_{0}\in\mathbb{N}$ will be chosen in the following sections and will depend on the value $b$. In particular, $n_{0}$ gets larger as $b$ is closer to $1$. We assume that the $Q$-Wiener process $z$ takes values $\mathbb{P}$-almost surely in the Sobolev space $H^{2+N_{1}}(\mathbb{T}^{3},\mathbb{R}^{3})$ with zero mean and divergence.
@@ 2.3. Preparation of the convex integration scheme
is larger than $T$ with probability greater than $\varkappa\in(0,1)$ given by the statement of ˜1.1. Notice that is always possible to find such $L$ since $H^{2+N_{1}}\subset C^{N_{1}}$ by Sobolev embedding. In this way, we have a $\mathbb{P}$-almost sure control over the size of $\|z_{t}\|_{N_{1}}$ on the time interval $t\in[0,\mathfrak{t}]$. However, since $z$ is temporally rough, we need to introduce a time mollification
@@
(2.12) $i_{q}:=\begin{cases}\lambda_{0}^{m_{5}}\lambda_{1}^{m_{6}},\,&q=0,\\ \lambda_{q}^{m_{1}}\lambda_{q+1}^{m_{2}}\delta_{q}^{m_{3}}\delta_{q+1}^{m_{4}},\,&q\geq 1,\end{cases}$
where $(m_{1},m_{2},m_{3},m_{4})=\left(1/2-\delta\right)^{-1}\left(0,0,-2,7/2\right)$ and $(m_{5},m_{6})=(7\alpha/(1-2\delta),-m_{5}).$ The choice of the asymmetric mollifier guarantees that $z_{q}$ remains adapted to the filtration $\{\mathbb{F}_{t}\}_{t\in\mathbb{R}}$, extended to negative times identically equal to $\mathbb{F}_{0}$. Hereafter, we will always adopt this convention when referring to the filtration $\{\mathbb{F}_{t}\}_{t\in\mathbb{R}}$.
2.4. Induction scheme
The main idea of [Is22] is to obtain a solution of Euler equations satisfying the local energy inequality as a limit of smooth solutions of approximating systems, which include a Reynolds stress tensor $R$ in the momentum equation and a flux current $\varphi$ in the local energy inequality, and boil down to the Euler system when $R=0$ and $\phi=0$. In this sense, the size of the errors $R$ and $\phi$ give an indication of “how far” a solution of the approximating system is from being an exact solution of the Euler system.
In this work, the approximating system to (2.5) is slightly more involved, since we need to keep track of the random process $z$ affecting the dynamics. Throughout this work we shall deal with probabilistically strong solutions, i.e. our objects will be defined on the same stochastic basis $(\Omega,\mathbb{F},(\mathbb{F}_{t}),\mathbb{P})$ supporting the Wiener process $z$.
Definition 2.1.
We point out that our approximating Euler-Reynolds flow will solve (2.13) (in an analytically strong sense) also for negative times $t<0$, as we work with one-sided time mollifications in order to preserve adaptedness to the filtration $\{\mathbb{F}_{t}\}_{t\in\mathbb{R}}$. We shall impose zero external noise for negative times, so that the process $(v,p,R,\phi,z)$ will be deterministic for negative times (and thus progressively measurable with respect to the filtration $\{\mathbb{F}_{t}\}_{t\in\mathbb{R}}$).
Lemma 2.2 (Iteration Lemma).
Let $\alpha\in(0,1/7)$ and $E\in C^{1}_{b}((-\infty,2T],\mathbb{R})$. Then there exist parameters $a,b$ as above, $C_{v},\overline{M}\in(1,\infty)$ and $n_{0}\in\mathbb{N}$ such that the following holds.
@@ 2.4. Induction scheme
(2.18) $\|v_{q+1}-v_{q}\|_{0}+\frac{1}{\lambda_{q+1}}\|v_{q+1}-v_{q}\|_{1}$ $\leq C_{v}\overline{M}\delta_{q+1}^{1/2}\lambda_{q}^{-\gamma/2},$
(2.19) $\|p_{q+1}-p_{q}\|_{0}+\frac{1}{\lambda_{q+1}}\|p_{q+1}-p_{q}\|_{1}$ $\leq\overline{M}\delta_{q+1}.$
We point out that, by definition of $\|\cdot\|_{-1}$ and $\|\cdot\|_{-2}$, estimates (2.15), (2.16) and (2.17) on the advective derivatives are void in these specific cases. Notice also that $\mathfrak{t}\leq 2T$ $\mathbb{P}$-almost surely by definition, so that the system (2.13) is meaningful on the time interval $(-\infty,\mathfrak{t}]$.
Since $\delta_{q}\to 0$ exponentially fast, the estimates above imply $\|R_{q}\|_{0},\|\phi_{q}\|_{0}\to 0$ as $q\to\infty$. Moreover, (2.18) and (2.19) on the increments and interpolation guarantee convergence $v_{q}\to v$ and $p_{q}\to p$ in a space of continuous-in-time, Hölder-in-space functions. In fact, given the previous lemma, we can prove ˜1.1 arguing as follows.
Proof of ˜1.1.
@@
(2.21) $\|R_{0}(t)\|_{N}$ $\leq L^{2}+|E(t)|\leq\lambda_{0}^{N-\gamma}(L^{2}+|E(t)|)\lambda_{0}^{\gamma},$
(2.22) $\|D_{t,0}R_{0}(t)\|_{N-2}$ $\leq i_{0}^{-1/2-\delta}L^{2}+L^{3}+|E^{\prime}(t)|\leq\lambda_{0}^{N-1-\gamma}(L^{3}+|E^{\prime}(t)|)\lambda_{0}^{3\gamma/2},$
where we have used that $i_{0}^{-1/2-\delta}\leq\lambda_{0}^{1+\gamma/2}$, as well as the bounds on the current
(2.23) $\|\phi_{0}(t)\|_{N}\leq L^{3}\leq L^{3}\lambda_{0}^{N-3\gamma/2}\lambda_{0}^{3\gamma/2},\quad\|D_{t,0}\phi_{0}(t)\|_{N-1}\leq i_{0}^{-1/2-\delta}L^{3}+L^{4}\leq L^{4}\lambda_{0}^{N-3\gamma/2}\lambda_{0}^{2\gamma},$
and the pressure
(2.24) $\|p_{0}(t)\|_{N}\leq L^{2}\leq L^{2}\lambda_{0}^{N},\quad\|D_{t,0}p_{0}(t)\|_{N-2}\leq i_{0}^{-1/2-\delta}L^{2}+L^{3}\leq L^{3}\lambda_{0}^{N-1}\lambda_{0}^{\gamma/2}.$

### v3

# Dissipative solutions to randomly forced 3D Euler equations

@@ Abstract.
The purpose of this work is twofold. First, we construct probabilistically strong solutions to the three-dimensional Euler equations perturbed by additive noise that are $\mathbb{P}$-almost surely continuous in time, Hölder in space, and satisfy the local energy inequality up to an arbitrarily large stopping time. Second, we prove several non-unique ergodicity results for the forced Euler equations with continuous-in-time external forcing. The solutions we construct are genuinely random and, almost surely, strictly dissipative and not steady states.
@@
Stochastic Euler equations, Local energy inequality, Convex integration, Ergodicity
@@ 1. Introduction
In the first part of this work, we consider the stochastic Euler equations on the three-dimensional torus $\mathbb{T}^{3}:=\mathbb{R}^{3}/2\pi\mathbb{Z}^{3}$ driven by additive noise
@@
Euler equations govern the evolution of an ideal fluid and constitute one of the most important models in fluid dynamics. The presence of a random external force in (SE) can help represent many phenomena perturbing the dynamics that are neglected by the deterministic equation, such as endogenous microscopic thermal effects or exogenous influences of the environment on the fluid. First investigations on (SE) date back to the works [MiVa00, Ki09], where local existence of smooth solutions is proved (see also [BeFl99] for other results in dimension two). Measure-valued solutions satisfying weak-strong uniqueness and a version of the energy inequality have been constructed in [BrMo21], building upon results by DiPerna and Majda [DPMa87] for the deterministic Euler equations $Q=0$. Whether solutions are unique or not remained an open problem until the work [HoZhZh22] by Hofmanová, Zhu, and Zhu, adapting for the first time the convex integration scheme developed by De Lellis and Székelyhidi Jr. in [DLSz09] to the stochastic case, and obtaining non-uniqueness in law for (SE). After [HoZhZh22] made clear that convex integration techniques can be applied to stochastic equations (let us also mention [ChFeFl21] in this regard), a sequence of papers has appeared [HoZhZh25, LuZh24, KiKo24, LuLuZh25] extending the works [DLS13, DLS14, BuDLIsSz15, BuDLIsSz16, Is18, BuDLSzVi19] on the deterministic Euler equations. In particular, for every $\vartheta\in(0,1/3)$, Lü, Lü, and Zhu proved in [LuLuZh25] existence and non-uniqueness of solutions of (SE) with regularity $u\in C([0,\infty),C^{\vartheta}(\mathbb{T}^{3},\mathbb{R}^{3})$ almost surely, which are strictly dissipative up to an arbitrarily large stopping time $\mathfrak{t}$ in the sense that $u$ satisfies almost surely the energy inequality
@@
The importance of a local form of energy inequality towards partial regularity results for suitable weak solutions of the Navier-Stokes equations is known since the works of Scheffer [Sc77] and Caffarelli, Kohn, and Nirenberg [CaKoNi82]. Similar results are available for the stochastic Navier-Stokes equations, see [Ro10, FlRo02]. The local energy balance (LEI) is a natural condition to impose also on the stochastic Euler equations (SE), since any physically relevant $L^{3}_{t,x}$ zero-viscosity limit of suitable solutions of the stochastic Navier-Stokes equations must inherit this property. It describes the local change in kinetic energy due to spatial roughness of $u$: indeed, the equality in (LEI) holds for smooth solutions $u$, and in general the inequality corresponds to no local creation of kinetic energy due to irregularities of $u$. We refer to the celebrated work of Duchon and Robert [DuRo00] for more details.
In the deterministic setting $Q=0$, existence and non-uniqueness of solutions to the system (SE)-(LEI) have been shown in the series of papers [DLSz10, Is22, DLK23, GiKwNo]. In the present work we provide similar results in the stochastic case $Q\neq 0$. More specifically, we will prove the following:
@@
Let $(\Omega,\mathbb{F},\{\mathbb{F}_{t}\}_{t\geq 0},\mathbb{P})$ be a stochastic basis with right-continuous and complete filtration supporting a sufficiently smooth-in-space $Q$-Wiener process $Q^{1/2}W$ with zero space average and null divergence. Let $T>0$, $\varkappa\in(0,1)$, $\alpha\in(0,1/7)$ and $E\in C^{1}_{loc}([0,\infty),\mathbb{R})$ be given. Then there exist two $\{\mathbb{F}_{t}\}_{t\geq 0}$-progressively measurable stochastic processes
(1.2) $u:\Omega\to C([0,\infty),C^{\alpha}(\mathbb{T}^{3},\mathbb{R}^{3})),\qquad p:\Omega\to C([0,\infty),C^{2\alpha}(\mathbb{T}^{3},\mathbb{R})),\qquad\mathbb{P}\mbox{-almost surely,}$
@@
In the second part of the paper, we move to addressing the problem of ergodicity associated with Euler equations. The so-called ergodic hypothesis postulates the existence of an equilibrium measure $\mu$ over some phase space $H$ of velocity fields $u$, such that the time average of observables $F(u_{t})$ over large time intervals tends to $\int_{H}Fd\mu$. This ergodic assumption is at the basis of several foundational studies of turbulence (see [FMRT01, Chapter IV] for a thorough discussion) but its validity lacks a rigorous mathematical justification. Therefore, in order to gain additional insight into the behaviour of fluids in a turbulent regime, it is of the utmost importance to consider ergodic measures $\mu$ concentrated on solutions to (SE). We point out that the presence of an external forcing introducing energy into the system (on average) is somehow necessary if one is interested in isolating statistically stationary solutions that are strictly dissipative.
Ideally, $\mu$ should be a suitable vanishing-viscosity limit of ergodic measures $\mu^{\nu}$ associated with the stochastically forced Navier-Stokes equations, cf. also the discussion at page 472 in [Ku04]. However, the recent work [HoZhZh25] has shown that (SE) admits non-unique ergodic invariant measures; Moreover, [HoZhZh25, Theorem 1.5] states that there exists $\vartheta>0$ with the property that every statistically stationary solution of (SE) with finite moments locally in $C_{t}H^{\vartheta}_{x}\cap C^{\vartheta}L^{2}_{x}$ is a vanishing-viscosity limit in law of statistically stationary solutions of stochastically forced Navier-Stokes equations. Therefore, vanishing viscosity alone is not a good selection criterion among statistically stationary solutions of (SE) (cf. also [BuVi19, Theorem 1.3] for a deterministic version of this statement).
Nonetheless, extending Theorem 1.1 to construct non-unique ergodic measures concentrated on solutions to (SE) satisfying the local energy inequality (LEI) is not straightforward. Let us only mention that one major issue is the unboundedness of the external forcing $Q^{1/2}dW$. On the other hand, we can construct non-unique ergodic measures $\mu$ on the path space
(1.3) $\mathscr{X}:=\{(u,p,g)\in C(\mathbb{R},C^{\alpha}(\mathbb{T}^{3},\mathbb{R}^{3}))\times C(\mathbb{R},C^{2\alpha}(\mathbb{T}^{3},\mathbb{R}))\times C(\mathbb{R},C^{2\alpha}(\mathbb{T}^{3},\mathbb{R}^{3})),\,\mathord{{\rm div}}(u)=\mathord{{\rm div}}(g)=0\},$
ergodicity being meant with respect to the shift operator on the path space [^1] Namely, $\mu(A)\in\{0,1\}$ for every shift invariant Borel set $S_{t}A=A\subset\mathscr{X}$, where $S_{t}(u,p,g)(s):=(u,p,g)(t+s)$. , that are concentrated on strictly dissipative solutions of the forced Euler equations
(FE) $\begin{cases}\partial_{t}u+\mathord{{\rm div}}(u\otimes u)+\nabla p=g,\\ \mathord{{\rm div}}(u)=0,\\ \frac{1}{2}\partial_{t}|u|^{2}+\mathord{{\rm div}}\left(\left(\frac{1}{2}|u|^{2}+p\right)u\right)dt-u\cdot g<0,\end{cases}$
with some external forcing $g$ having a given law, and such that $u$ is genuinely random and almost surely not a steady-state. More precisely, we have the following:
Theorem 1.2.
There exist two distinct ergodic measures $\mu_{1},\mu_{2}$ on the path space $\mathscr{X}$, concentrated on solutions of (FE), such that for $i\in\{1,2\}$ the process $u$ is not a steady state
(1.4) $\sup_{s,t\in\mathbb{R}}\|u(t)-u(s)\|_{C_{x}}>0,\quad\mu_{i}\,\mbox{-\,almost surely},$
and is genuinely random
(1.5) $\inf_{t\in\mathbb{R}}\mathbb{E}^{\mu_{i}}\left[\|u(t)-\mathbb{E}^{\mu_{i}}[u(t)]\|^{2}_{C_{x}}\right]>0.$
Moreover, we can construct $\mu_{1},\mu_{2}$ such that the laws of $g$ under $\mu_{1}$ and $\mu_{2}$ coincide, but the laws of $u$ under $\mu_{1}$ and $\mu_{2}$ differ. In particular, for strictly dissipative ergodic measures associated with (FE)-(1.4)-(1.5) one can not uniquely define a map $\mathscr{L}_{\mu}(g)\mapsto\mathscr{L}_{\mu}(u)$.
Remark 1.3.
The validity of (1.4)-(1.5) is a well-known open problem in limiting procedures à la Kuksin [Ku04], see for example the discussion in [BeCZGH16, Section 2.3].
What distinguishes the two ergodic measures $\mu_{1},\mu_{2}$ in the previous theorem is the typical size of the process $u$, which also entails that the amount of kinetic energy dissipated under $\mu_{1}$ and $\mu_{2}$ differs. If we don’t prescribe the law of the external forcing $g$ then we can fix the right-hand side in the third line of (FE). In addition, similarly to [HoZhZh25, Theorems 6.1 - 6.3] we can almost arbitrarily prescribe the law of $u$ up to a small error.
Theorem 1.4.
For every $\iota\in(0,1/8)$, constant $E^{\prime}>0$ sufficiently small and every shift-invariant law $\nu$ on the space $\mathscr{Z}:=\{Z\in C(\mathbb{R},C^{1}(\mathbb{T}^{3},\mathbb{R}^{3}),\,\mathord{{\rm div}}(Z)=0,\,\inf_{t\in\mathbb{R}}\|Z(t+1)-Z(t)\|_{C_{x}}>\frac{1}{2}\|Z\|_{C_{t,x}}\}$ satisfying the assumptions of Lemma 2.3, there exists an ergodic measure $\mu$} on the path space $\mathscr{X}\times\mathscr{Z}$, concentrated on solutions of (FE) with dissipation equal to $E^{\prime}$, such that (1.4) and (1.5) hold, $\mathscr{L}_{\mu}(Z)=\nu$, and
(1.6) $\sup_{t\in\mathbb{R}}\|u(t)-Z(t)\|_{C_{x}}<\iota\inf_{t\in\mathbb{R}}\|Z(t)\|_{C_{x}},\quad\mu\,\mbox{-\,almost surely}.$
The paper is organized as follows. In section 2 we give some preliminaries and introduce the convex integration schemes that we will use to prove our theorems. Iterative estimates for (SE) are contained in section 3 (for the velocity field), section 4 (for the Reynolds stress and pressure), and in section 5 (for the current). These results are also at the basis of the iteration lemma needed to construct solutions of (FE), and are complemented by the results in section 6 that are specific to this case. Some auxiliary results are collected in the Appendix.
@@ Acknowledgements
The authors are sincerely grateful to Eliseo Luongo and Marco Romito for many interesting discussions. UP has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 949981) and the Swiss National Science Foundation under the SNSF Ambizione grant No. 233216.
@@
2. Preliminaries and main iterative propositions
2.1. Notation
Hölder spaces $C^{\gamma}(\mathcal{O},\mathbb{R}^{m})$, $\gamma\geq 0$ are defined according to the following notion of norms. Let $N\in\mathbb{N}$ and $\delta\in(0,1)$, and let us introduce the following Hölder seminorms:
Then, the Hölder norms are defined as
Furthermore, in order to simplify the statement of many propositions here contained, we adopt the notational convention $\|f\|_{M}\equiv 0$ whenever the index $M<0$. For $\gamma\geq 0$ we denote $C^{\gamma}_{x}:=C^{\gamma}(\mathcal{O},\mathbb{R}^{m})$, with $\mathcal{O}$ and $m$ depending on the context, and similarly for other functional spaces. We use the symbol $c^{\gamma}_{x}$ to denote the closure of smooth functions with respect to the topology of $C^{\gamma}_{x}$.
2.2. Preparation of the convex integration scheme
The proofs of Theorem 1.1, Theorem 1.2 and Theorem 1.4 rely on a convex integration scheme, mostly inspired by the works [Is22, DLK23, LuLuZh25]. Let us introduce various parameters that will allow to estimate all the objects appearing in the upcoming sections. Define for $q\in\mathbb{N}$:
where $a>2$ will be very large, $b>1$ will be chosen close to 1, $\gamma:=(b-1)^{2}$, and $\alpha\in(0,1/7)$ indicates the Hölder regularity of the solution identified by the iterative procedure.
2.3. Stochastic Euler Equations: Induction scheme
In this subsection, we intend to establish Theorem 1.1 by making use of an Iteration Lemma (see Lemma 2.1), whose proof will be given in the following sections.
@@ 2.3. Stochastic Euler Equations: Induction scheme
(2.7) $\partial_{t}\frac{|v|^{2}}{2}+v\cdot\nabla p+v\cdot\mathord{{\rm div}}\big((v+z)\otimes(v+z)\big)=-E^{\prime}.$
@@
We aim at building solutions of (SE)-(LEE) by applying the convex integration scheme to (2.8). In particular, following [Is22], we obtain a solution of (2.8) as a limit of smooth solutions of approximating systems, which include a Reynolds stress tensor $R$ in the momentum equation and a flux current $\phi$ in the local energy inequality, and boil down to the (2.8) when $R=0$ and $\phi=0$. In this sense, the size of the errors $R$ and $\phi$ give an indication of “how far” a solution of the approximating system is from being an exact solution.
The approximating system to (2.8) is slightly more involved than the one introduced at [DLK23], since we need to keep track of the stochastic process $z$ affecting the dynamics. Whenever we work on Stochastic Euler equations, we shall deal with probabilistically strong solutions, i.e. our objects will be defined on the same stochastic basis $(\Omega,\mathbb{F},\{\mathbb{F}_{t}\}_{t\in\mathbb{R}},\mathbb{P})$ supporting the Wiener process $z$.
Definition 2.1.
We point out that our approximating Euler-Reynolds flow will solve (2.9) (in an analytically strong sense) also for negative times $t<0$, as we work with one-sided time mollifications in order to preserve adaptedness to the filtration $\{\mathbb{F}_{t}\}_{t\in\mathbb{R}}$. We shall impose zero external noise for negative times, so that the process $(v,p,R,\phi,z)$ will be $\mathbb{F}_{0}$-adapted for negative times (and thus progressively measurable with respect to the filtration $\{\mathbb{F}_{t}\}_{t\in\mathbb{R}}$).
Before reporting the Iteration Lemma, let us introduce various assumptions on $z$ and define $z_{q}$. Let $N_{1}:=n_{0}+5$, where the parameter $n_{0}\in\mathbb{N}$ will be chosen in the following sections and will depend on the value $b$. In particular, $n_{0}$ gets larger as $b$ is closer to $1$. We assume that the $Q$-Wiener process $z$ takes values $\mathbb{P}$-almost surely in the Sobolev space $H^{2+N_{1}}(\mathbb{T}^{3},\mathbb{R}^{3})$ with zero mean and divergence.
@@
is larger than $T$ with probability greater than $\varkappa\in(0,1)$ given by the statement of Theorem 1.1. Notice that is always possible to find such $L$ since $H^{2+N_{1}}_{x}\subset C^{N_{1}}_{x}$ by Sobolev embedding. In this way, we have a $\mathbb{P}$-almost sure control over the size of $\|z_{t}\|_{N_{1}}$ on the time interval $t\in[0,\mathfrak{t}]$. However, since $z$ is temporally rough, we build $z_{q}$ by introducing a time mollification
@@
(2.13) $i_{q}^{1/2-\delta}:=\begin{cases}\lambda_{0}^{7\alpha/2}\lambda_{1}^{-7\alpha/2},\,&q=0,\\ \delta_{q}^{-2}\delta_{q+1}^{7/2},\,&q\geq 1.\end{cases}$
The choice of the asymmetric mollifier guarantees that $z_{q}$ remains adapted to the filtration $\{\mathbb{F}_{t}\}_{t\in\mathbb{R}}$, extended to negative times identically equal to $\mathbb{F}_{0}$. Hereafter, we will always adopt this convention when referring to the filtration $\{\mathbb{F}_{t}\}_{t\in\mathbb{R}}$.
Lemma 2.1.
Let $\alpha\in(0,1/7)$ and $E\in C^{1}_{b}((-\infty,2T],\mathbb{R})$. Then there exist parameters $a,b$ as above, $n_{0}\in\mathbb{N}$ depending only on $b$, and $C_{v},\overline{M}\in(1,\infty)$ such that the following holds.
@@
(2.18) $\|v_{q+1}(t)-v_{q}(t)\|_{0}+\frac{1}{\lambda_{q+1}}\|v_{q+1}(t)-v_{q}(t)\|_{1}$ $\leq C_{v}\overline{M}\delta_{q+1}^{1/2}\lambda_{q}^{-\gamma/2},$
(2.19) $\|p_{q+1}(t)-p_{q}(t)\|_{0}+\frac{1}{\lambda_{q+1}}\|p_{q+1}(t)-p_{q}(t)\|_{1}$ $\leq\overline{M}\delta_{q+1}.$
We point out that, by definition of $\|\cdot\|_{-1}$ and $\|\cdot\|_{-2}$, estimates (2.15), (2.16) and (2.17) on the advective derivatives are void in these specific cases. Notice also that $\mathfrak{t}\leq 2T$ by definition, so that the system (2.9) is meaningful on the time interval $(-\infty,\mathfrak{t}]$.
Since $\delta_{q}\to 0$ exponentially fast, the estimates above imply $\|R_{q}\|_{0},\|\phi_{q}\|_{0}\to 0$ as $q\to\infty$. Moreover, (2.18) and (2.19) on the increments and interpolation guarantee convergence $v_{q}\to v$ and $p_{q}\to p$ in a space of continuous-in-time, Hölder-in-space functions. In fact, given the previous lemma, we can prove Theorem 1.1 arguing as follows.
Proof of Theorem 1.1.
@@
(2.21) $\|R_{0}(t)\|_{N}$ $\leq|E(t)|+\sum_{k=0}^{N}\binom{N}{k}\|z_{0}\|_{k}\|z_{0}\|_{N-k}$
(2.22) $\leq|E(t)|+(N+1)2^{N}L^{2}\leq\lambda_{0}^{N-\gamma}(L^{2}+|E(t)|)\lambda_{0}^{\gamma},$
(2.23) $\|D_{t,0}R_{0}(t)\|_{N-2}$ $\leq|E^{\prime}(t)|+(N-1)2^{N-2}i_{0}^{-1/2-\delta}L^{2}+\sum_{k+h=0}^{N-2}\binom{N-2}{k,h}\|z_{0}\|_{k}\|z_{0}\|_{h}\|z_{0}\|_{N-1-k-h}$
(2.24) $\leq|E^{\prime}(t)|+(N-1)2^{N-2}i_{0}^{-1/2-\delta}L^{2}+(N-1)^{2}3^{N-2}L^{3}\leq\lambda_{0}^{N-1-\gamma}(L^{3}+|E^{\prime}(t)|)\lambda_{0}^{3\gamma/2},$
where we have used that $i_{0}^{-1/2-\delta}\leq\lambda_{0}^{1+\gamma/2}$ and we have taken $a$ sufficiently large, as well as the bounds on the current
(2.25) $\|\phi_{0}(t)\|_{N}$ $\leq(N+1)^{2}3^{N}L^{3}\leq L^{3}\lambda_{0}^{N-3\gamma/2}\lambda_{0}^{3\gamma/2},$
(2.26) $\|D_{t,0}\phi_{0}(t)\|_{N-1}$ $\leq N^{2}3^{N-1}i_{0}^{-1/2-\delta}L^{3}+N^{3}4^{N-1}L^{4}\leq L^{4}\lambda_{0}^{N-3\gamma/2}\lambda_{0}^{2\gamma},$
and on the pressure
(2.27) $\|p_{0}(t)\|_{N}$ $\leq(N+1)2^{N}L^{2}\leq L^{2}\lambda_{0}^{N},$
(2.28) $\|D_{t,0}p_{0}(t)\|_{N-2}$ $\leq(N-1)2^{N-2}i_{0}^{-1/2-\delta}L^{2}+(N-1)^{2}3^{N-2}L^{3}\leq L^{3}\lambda_{0}^{N-1}\lambda_{0}^{\gamma/2}.$